Heather reeived grades of 80, 75, and 86 on three algebra tests. What must her grade be on the next test if her average for the four tests is to be at least 82? How do I solve this problem?
Answer by nerdybill(7384)

Heather reeived grades of 80, 75, and 86 on three algebra tests. What must her grade be on the next test if her average for the four tests is to be at least 82? How do I solve this problem?
.
Let x = score on fourth test
then
(80+75+86+x)/4 = 82
(241+x)/4 = 82
241+x = 82*4
241+x = 328
x = 87
.
That is Heather's score must be >= 87